Question2 1.5 points Save Answen Find the output of the following python code a-10 b15 a-a+b print b Path:p Words:0
Answer #1

Answer:

a=10 /* a is initialized by 10 */
b=15 /* b is initialized by 15 */
a=a+b /* The contents of a and b are added and the result is stored in a (a=10+15==>a=25) */
print b /* print b is invalid syntax. Instead of writing print b, we shold write it as print (b) */

The program shows syntax error at print b (The syntax error is Invalid syntax)

if the question is as follows

a=10
b=15
a=a+b
print (b) /* prints b value (that is 15) */